# Differentially-Private Distributed Model Predictive Control of Linear Discrete-Time Systems with Global Constraints

### v1

@@ 4             Every subsystem $i$ sends $\lambda_{i}^{k}$ to subsystem $j\in\mathcal{N}_{i}$;
1 At time instant $t$, each subsystem $i$ measures its state $x_{i}(t)$;
2Each subsystem $i$ obtains $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 1 with $x_{i}(t)$;
@@ 1 At time instant $t$, each subsystem $i$ measures its state $x_{i}(t)$;
Wait for the next sampling instant; let $t=t+1$ and go to step 1.
@@
If Assumptions 2 and 3 hold and if the step-size $\gamma^{k}$ satisfies the conditions $\sum_{k=0}^{\infty}\gamma^{k}=\infty$, $\sum_{k=0}^{\infty}(\gamma^{k})^{2}<\infty$, then Algorithm 1 guarantees the convergence of the sequence $\{\tilde{{\bm{u}}}_{i}^{k}\}$, i.e., $\lim_{k\rightarrow\infty}\|\tilde{{\bm{u}}}_{i}^{k}-\tilde{{\bm{u}}}_{i}^{*}\|$. Note that the objective function of problem (4) is strictly convex, and thus the asymptotic primal convergence can be established without resorting to local averaging mechanism (see Section 3.4 in Notarstefano et al. (2019) for more details). In addition, if the iteration number $\bar{k}$ is selected sufficiently large to meet specific criteria (Su et al., 2022; Wang and Ong, 2017), then Algorithm 2 ensures the feasibility and stability of the considered MPC problem.
@@
2.3 On Differential Privacy
In this work, DP is used to characterize and quantify the achieved privacy level of distributed optimization algorithms. Given the continual exchange of information among subsystems in iterative optimization algorithms, the notion of $\epsilon$-DP for continuous bit streams (Dwork et al., 2010) is adopted. Drawing inspiration from the distributed optimization framework proposed by Huang et al. (2015), we represent the DMPC problem in (4) by four parameters $(\mathcal{J},\tilde{\mathcal{U}},\mathcal{G},L)$ to facilitate DP analysis. Specifically, $\mathcal{J}:=\{J_{1},\,\cdots,J_{M}\}$ denotes the set of objective functions for individual subsystems, $\tilde{\mathcal{U}}:=\{\tilde{\mathcal{U}_{1}},\cdots,\tilde{\mathcal{U}_{M}}\}$ is the domain of optimization variables, $\mathcal{G}:=\{g_{1},\cdots,g_{M}\}$ represents the set of constraint functions for individual subsystems, and $L$ is the weight matrix describing the communication network. The adjacency between two optimization problems is defined as follows:
@@ 2.3 On Differential Privacy
Two optimization problems $\mathcal{P}=(\mathcal{J},\tilde{\mathcal{U}},\mathcal{G},L)$ and $\mathcal{P}^{{}^{\prime}}=(\mathcal{J}^{{}^{\prime}},\tilde{\mathcal{U}}^{{}^{\prime}},\mathcal{G}^{{}^{\prime}},L^{{}^{\prime}})$ are adjacent if the following conditions hold:
  • 1.
    the objective functions, the domains of optimization variables, and the interaction weight matrices are identical, i.e., $\mathcal{J}=\mathcal{J}^{{}^{\prime}}$, $\tilde{\mathcal{U}}=\tilde{\mathcal{U}}^{{}^{\prime}}$, and $L=L^{{}^{\prime}}$;
  • 2.
    there exists an $i\in\mathbb{Z}_{1}^{M}$ such that $g_{i}\neq g_{i}^{{}^{\prime}}$ but $g_{j}=g_{j}^{{}^{\prime}}$ for all $j\in\mathbb{Z}_{1}^{M},\,j\neq i$;
  • 3.
    $g_{i}$ and $g_{i}^{{}^{\prime}}$, which are not the same, have similar behaviors around $\theta^{*}$, the solution of $\mathcal{P}$. More specifically, there exists some $\delta>0$ such that for all $\bm{u}_{i}$ and $\bm{u}_{i}^{{}^{\prime}}$ in $B_{\delta}(\theta^{*}):=\{\bm{v}:\bm{v}\in\mathbb{R}^{Nm_{i}},\|\bm{v}-\theta^{*}\|<\delta\}$, we have $g_{i}(\bm{u}_{i})=g_{i}^{{}^{\prime}}(\bm{u}_{i}^{{}^{\prime}})$.
In the context of a distributed optimization algorithm, we denote an execution of such an algorithm as $\mathcal{A}$, represented by a sequence of the iteration variable $\vartheta$, i.e., $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\cdots\}$. We consider adversaries that can get access to all communicated messages among the subsystems. Hence, under an execution $\mathcal{A}$, the observation of adversaries is the sequence of shared messages, which is denoted as $\mathcal{O}$. For a given distributed optimization problem $\mathcal{P}$ and an initial state $\vartheta^{0}$, we define the observation mapping as $\mathcal{R}_{\mathcal{P},\vartheta^{0}}(\mathcal{A}):=\mathcal{O}$. Given a distributed optimization problem $\mathcal{P}$, an observation sequence $\mathcal{O}$, and an initial state $\vartheta^{0}$, $\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O})$ denotes the set of executions $\mathcal{A}$ capable of generating the observation $\mathcal{O}$.
Definition 2 ($\epsilon$-differential privacy, Huang et al. (2015)).
For a given $\epsilon>0$, an iterative distributed algorithm is $\epsilon$-differentially private if for any two adjacent $\mathcal{P}$ and $\mathcal{P}^{\prime}$, any set of observation sequences $\mathcal{O}_{s}\subseteq\mathbb{O}$ (with $\mathbb{O}$ denoting the set of all possible observation sequences), and any initial state ${\vartheta}^{0}$, the following relationship always holds
@@
with the probability $\mathbb{P}$ taken over the randomness over iteration processes.
The definition of $\epsilon$-DP guarantees that adversaries, with access to all shared messages, cannot gain knowledge about any participating subsystem’s sensitive information. It can be seen that a smaller $\epsilon$ means a higher level of privacy protection.
3 Differentially-Private Distributed Dual-Gradient Algorithm
3.1 Algorithm Description
@@
In contrast to Algorithm 1, where each subsystem directly sends $\lambda_{i}^{k}$ to its neighbors, Algorithm 3 incorporates DP noise $\zeta_{i}^{k}$ into $\lambda_{i}^{k}$ and shares the perturbed signal $\hat{\lambda}_{i}^{k}:=\lambda_{i}^{k}+\zeta_{i}^{k}$ among the communication network. Therefore, the information available to potential adversaries is the sequence $\{\hat{\lambda}_{i}^{k}\}$. Due to the randomness of DP noise, it is impossible for the adversary to extract useful information from $\{\hat{\lambda}_{i}^{k}\}$ with significant probability. Furthermore, it should be noted that directly integrating persistent DP noise into existing optimization algorithms will compromise the convergence accuracy. To address this issue, we utilize findings from Wang (2023); Wang and Nedić (2024) to design a weakening factor. As shown in (18), the weakening factor, denoted as $\chi^{k}$, is applied on the interaction terms ($L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})$). The fundamental principle behind incorporating this weakening factor is to gradually eliminate the impact of DP noise on convergence accuracy.
@@
$\lambda_{i}^{k+1}=\Pi_{\mathbb{R}^{Np}_{+}}\left[\lambda_{i}^{k}+\chi^{k}\sum_{j\in\mathcal{N}_{i}}L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})+\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right];$ (18)
@@
To facilitate the convergence and privacy analysis, we make the following assumption on the DP noise:
@@
For every $i\in\mathbb{Z}_{1}^{M}$ and every $k$, conditional on $\lambda_{i}^{k}$, the DP noise $\zeta_{i}^{k}$ satisfies $\mathbb{E}\left[\zeta_{i}^{k}\mid\lambda_{i}^{k}\right]=0$ and $\mathbb{E}\left[\|\zeta_{i}^{k}\|^{2}\mid\lambda_{i}^{k}\right]=(\sigma_{i}^{k})^{2}$ for all $k\geq 0$, and
@@
where $\{\chi^{k}\}$ is the weakening factor sequence from Algorithm 3. The initial random vectors satisfy $\mathbb{E}\left[\|\lambda_{i}^{0}\|^{2}\right]<\infty$, $\forall i\in\mathbb{Z}_{i}^{M}$.
Considering Assumption 4, we use the Laplace noise mechanism to generate $\zeta_{i}^{k}$, i.e., we add Laplace noise to all shared messages. More specifically, given a constant $\nu>0$, we use ${\rm Lap}(\nu)$ to denote a Laplace distribution of a scalar random variable with the probability density function being $\rho\rightarrow\frac{1}{2\nu}e^{-\frac{|\rho|}{\nu}}$. At each iteration $k$, all elements of $\zeta_{i}^{k}$ are drawn independently from Laplace distribution ${\rm Lap}(\nu^{k})$, where $\nu^{k}>0$. One can verify that the mean and variance of ${\rm Lap}(\nu^{k})$ is zero and $2(\nu^{k})^{2}$, respectively. Therefore, $\zeta_{i}^{k}$ satisfies $\mathbb{E}\left[\zeta_{i}^{k}\mid\lambda_{i}^{k}\right]=0$ and $\mathbb{E}\left[\|\zeta_{i}^{k}\|^{2}\mid\lambda_{i}^{k}\right]=(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$.
@@ 3.2 Convergence Analysis
According to Assumptions 3, 4, (21), and (22), we can follow the same line of reasoning as that of Theorem 1 in Wang and Nedić (2024) to obtain the results.
We also need the following lemma for convergence analysis:
Lemma 1 (Lemma 11, Polyak (1987)).
@@
$\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*}),\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k}\},\lambda^{*})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$
@@
Based on Lemma 1 in Nedić et al. (2010) and the update law of $\lambda_{i}^{k}$ in (18), it can be obtained that for any $\lambda\in\mathbb{R}_{+}^{Np}$,
@@
$\qquad\left.+2\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right)^{\top}\left(\chi^{k}\xi_{i}^{k}\right)+2\left(\bar{\lambda}^{k}-\lambda\right)^{\top}\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)\right.$
@@
$d^{k}=(\chi^{k})^{2}\sum_{i=1}^{M}\sum_{j\in\mathcal{N}_{i}}(L_{ij}\sigma_{j}^{k})^{2}+M(\gamma^{k})^{2}C_{g}^{2}+6C_{g}\gamma^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|.$ (30)
@@
Since $\gamma^{k}$ is non-summable, we have that ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ and ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})$ converge to zero almost surely.
@@
3.3 Privacy Analysis
Based on the adjacency concept delineated in Definition 1, we can establish two adjacent distributed optimization problems, denoted as $\mathcal{P}$ and $\mathcal{P}^{\prime}$. The only difference between these two problems is a single signal (represent it as $g_{i}$ in $\mathcal{P}$ and $g^{\prime}_{i}$ in $\mathcal{P}^{\prime}$ without loss of generality). In accordance with the third condition of Definition 1, signals $g_{i}$ and $g^{{}^{\prime}}_{i}$ are required to exhibit similar behaviors around the optimal solution, i.e., $g_{i}$ and $g^{{}^{\prime}}_{i}$ should converge to each other if the algorithm can ensure convergence to the optimal solution. Therefore, based on the proven convergence in Theorem 2, we can formalize this condition by stipulating the existence of a constant $C>0$ such that

### v2

@@ 4             Every subsystem $i$ sends $\lambda_{i}^{k}$ to subsystem $j\in\mathcal{N}_{i}$;
1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
2Every subsystem $i$ computes $\tilde{{\bm{u}}}_{i}^{\bar{k}}$ by following Algorithm 1 with $x_{i}(t)$;
@@ 1 At time instant $t$, every subsystem $i$ measures its state $x_{i}(t)$;
Wait for the next time instant; let $t=t+1$ and go to step 1.
@@
If Assumptions 2 and 3 hold and if the step-size $\gamma^{k}$ satisfies the conditions $\sum_{k=0}^{\infty}\gamma^{k}=\infty$, $\sum_{k=0}^{\infty}(\gamma^{k})^{2}<\infty$, then Algorithm 1 guarantees the convergence of the sequence $\{\tilde{{\bm{u}}}_{i}^{k}\}$, i.e., $\lim_{k\rightarrow\infty}\|\tilde{{\bm{u}}}_{i}^{k}-\tilde{{\bm{u}}}_{i}^{*}\|$. Note that the objective function of problem (4) is strictly convex, and thus the asymptotic primal convergence can be established without resorting to local averaging mechanism (see Section 3.4 in [28] for more details). In addition, if the iteration number $\bar{k}$ is selected sufficiently large to meet specific criteria [29, 11], then Algorithm 2 ensures the feasibility and stability of the considered MPC problem.
@@
II-C On Differential Privacy
In this work, DP is used to characterize and quantify the achieved privacy level of distributed optimization algorithms. Given the continual exchange of information among subsystems in iterative optimization algorithms, the notion of $\epsilon$-DP for continuous bit streams [30] is adopted. Drawing inspiration from the distributed optimization framework proposed by [22], we represent the DMPC problem in (4) by four parameters $(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ to facilitate DP analysis. Specifically, $L$ is the interaction weight matrix describing the communication network, $\mathcal{J}:=\{J_{1},\,\cdots,J_{M}\}$ denotes the set of objective functions for individual subsystems, $\tilde{\mathcal{U}}:=\{\tilde{\mathcal{U}_{1}},\cdots,\tilde{\mathcal{U}_{M}}\}$ is the domain of optimization variables, and $\mathcal{G}:=\{g_{1},\cdots,g_{M}\}$ represents the set of constraint functions for individual subsystems. The adjacency between two optimization problems is defined as follows:
@@ II-C On Differential Privacy
Two distributed optimization problems $\mathcal{P}=(L,\mathcal{J},\tilde{\mathcal{U}},\mathcal{G})$ and $\mathcal{P}^{{}^{\prime}}=(L^{{}^{\prime}},\mathcal{J}^{{}^{\prime}},\tilde{\mathcal{U}}^{{}^{\prime}},\mathcal{G}^{{}^{\prime}})$ are adjacent if they satisfy the following conditions:
  • the interaction weight matrices, the objective functions, and the domains of optimization variables are identical, i.e., $L=L^{{}^{\prime}}$, $\mathcal{J}=\mathcal{J}^{{}^{\prime}}$, and $\tilde{\mathcal{U}}=\tilde{\mathcal{U}}^{{}^{\prime}}$;
  • there exists an $i\in\mathbb{Z}_{1}^{M}$ such that $g_{i}\neq g_{i}^{{}^{\prime}}$, and $g_{j}=g_{j}^{{}^{\prime}}$ for all $j\in\mathbb{Z}_{1}^{M},\,j\neq i$;
  • $g_{i}$ and $g_{i}^{{}^{\prime}}$, while different, exhibit similar behaviors near $\theta^{*}$, where $\theta^{*}$ is the solution of $\mathcal{P}$. More precisely, there exists a $\delta>0$ such that for all $\bm{u}_{i}$ and $\bm{u}_{i}^{{}^{\prime}}$ within the domain $B_{\delta}(\theta^{*}):=\{\bm{v}:\bm{v}\in\mathbb{R}^{Nm_{i}},\|\bm{v}-\theta^{*}\|<\delta\}$, $g_{i}(\bm{u}_{i})=g_{i}^{{}^{\prime}}(\bm{u}_{i}^{{}^{\prime}})$ holds.
We denote the execution of a distributed optimization algorithm as $\mathcal{A}$, represented by a sequence of the iteration variable $\vartheta$, i.e., $\mathcal{A}=\{\vartheta^{0},\vartheta^{1},\cdots\}$. We assume that adversaries have access to all communicated messages among the subsystems. Hence, under an execution $\mathcal{A}$, the adversaries’ observation is the sequence of communicated messages, denoted as $\mathcal{O}$. Let $\mathbb{O}$ represent the set of all possible observation sequences. For a distributed optimization problem $\mathcal{P}$ with an initial state $\vartheta^{0}$, the observation mapping is defined as $\mathcal{R}_{\mathcal{P},\vartheta^{0}}(\mathcal{A}):=\mathcal{O}$. Furthermore, for a distributed optimization problem $\mathcal{P}$, an initial state $\vartheta^{0}$, and an observation sequence $\mathcal{O}$, $\mathcal{R}_{\mathcal{P},\vartheta^{0}}^{-1}(\mathcal{O})$ denotes the set of executions $\mathcal{A}$ capable of generating the observation $\mathcal{O}$.
Definition 2 ($\epsilon$-differential privacy, [22]).
For a given $\epsilon>0$, an iterative distributed algorithm ensures $\epsilon$-differential privacy if for any two adjacent optimization problems $\mathcal{P}$ and $\mathcal{P}^{\prime}$, any initial state ${\vartheta}^{0}$, and any set of observation sequences $\mathcal{O}_{s}\subseteq\mathbb{O}$, the following relationship always holds:
@@
with the probability $\mathbb{P}$ taken over the randomness of iteration processes.
The definition of $\epsilon$-DP guarantees that adversaries, with access to all communicated information, cannot infer knowledge about any participating subsystem’s sensitive information. It can be found that a smaller $\epsilon$ indicates a better extent of privacy preservation.
III Differentially-Private Distributed Dual-Gradient Algorithm
III-A Algorithm Description
@@
In contrast to Algorithm 1, where each subsystem directly sends $\lambda_{i}^{k}$ to its neighbors, Algorithm 3 incorporates DP noise $\zeta_{i}^{k}$ into $\lambda_{i}^{k}$ and shares the perturbed signal $\hat{\lambda}_{i}^{k}:=\lambda_{i}^{k}+\zeta_{i}^{k}$ among the communication network. Therefore, the information available to potential adversaries is the sequence $\{\hat{\lambda}_{i}^{k}\}$. Due to the randomness of DP noise, it is impossible for the adversary to extract useful information from $\{\hat{\lambda}_{i}^{k}\}$ with significant probability. Furthermore, it should be noted that directly integrating persistent DP noise into existing optimization algorithms will compromise the convergence accuracy. To address this issue, we utilize findings from [25, 26] to design a weakening factor. As shown in (18), the weakening factor, denoted as $\chi^{k}$, is applied on the interaction terms ($L_{ij}(\hat{\lambda}_{j}^{k}-\lambda_{i}^{k})$). The fundamental principle behind incorporating this weakening factor is to gradually eliminate the impact of DP noise on convergence accuracy.
@@
[…]
@@
To facilitate the convergence and privacy analysis, the following DP noise assumption is introduced:
@@
For every $k$ and every $i\in\mathbb{Z}_{1}^{M}$, conditional on $\lambda_{i}^{k}$, the DP noise $\zeta_{i}^{k}$ satisfies $\mathbb{E}\left[\zeta_{i}^{k}\mid\lambda_{i}^{k}\right]=0$ and $\mathbb{E}\left[\|\zeta_{i}^{k}\|^{2}\mid\lambda_{i}^{k}\right]=(\sigma_{i}^{k})^{2}$ for all $k\geq 0$, and
@@
where $\{\chi^{k}\}$ is the weakening factor sequence from Algorithm 3.
Considering Assumption 4, we use the Laplace noise mechanism to generate $\zeta_{i}^{k}$ and then add it to all shared messages. More specifically, given a constant $\nu>0$, let ${\rm Lap}(\nu)$ represent a Laplace distribution of a scalar random variable, and $\rho\rightarrow\frac{1}{2\nu}e^{-\frac{|\rho|}{\nu}}$ be the corresponding probability density function. At each iteration $k$, every element of $\zeta_{i}^{k}$ is independently sampled from Laplace distribution ${\rm Lap}(\nu^{k})$, where $\nu^{k}>0$. One can verify that the mean and variance of ${\rm Lap}(\nu^{k})$ is zero and $2(\nu^{k})^{2}$, respectively. Therefore, $\zeta_{i}^{k}$ satisfies $\mathbb{E}\left[\zeta_{i}^{k}\mid\lambda_{i}^{k}\right]=0$ and $\mathbb{E}\left[\|\zeta_{i}^{k}\|^{2}\mid\lambda_{i}^{k}\right]=(\sigma_{i}^{k})^{2}=2(\nu^{k})^{2}$.
@@ III-B Convergence Analysis
According to Assumptions 3, 4, (21), and (22), we can follow the same line of reasoning as that of Theorem 1 in [26] to obtain the results. ∎
The following lemma is also required for convergence analysis:
Lemma 1 (Lemma 11, [31]).
@@
$\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*}),$
$\lim_{k\rightarrow\infty}\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k}\},\lambda^{*})=\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$
@@
Based on Lemma 1 in [32] and the update law of $\lambda_{i}^{k}$ in (18), it can be obtained that for any $\lambda\in\mathbb{R}_{+}^{Np}$,
@@
$\qquad\left.+2\left(\sum_{j\in\mathcal{N}_{i}\cup\{i\}}w_{ij}^{k}\lambda_{j}^{k}-\lambda\right)^{\top}\left(\chi^{k}\xi_{i}^{k}\right)\right.$
$\qquad\left.+2\left(\bar{\lambda}^{k}-\lambda\right)^{\top}\left(\gamma^{k}g_{i}(\tilde{{\bm{u}}}_{i}^{k+1})\right)\right.$
@@
$d^{k}=$ $(\chi^{k})^{2}\sum_{i=1}^{M}\sum_{j\in\mathcal{N}_{i}}(L_{ij}\sigma_{j}^{k})^{2}+M(\gamma^{k})^{2}C_{g}^{2}$ (30)
$+6C_{g}\gamma^{k}\sum_{i=1}^{M}\left\|\lambda_{i}^{k}-\bar{\lambda}^{k}\right\|.$
@@
Since $\gamma^{k}$ is non-summable, we have that ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\bar{\lambda}^{k})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})$ and ($\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{*}\},\lambda^{*})-\mathcal{L}(\{\tilde{{\bm{u}}}_{i}^{k+1}\},\lambda^{*})$ converge to zero almost surely. ∎
@@
III-C Privacy Analysis
Based on the adjacency concept delineated in Definition 1, we can establish two adjacent distributed optimization problems, denoted as $\mathcal{P}$ and $\mathcal{P}^{\prime}$. There is only one signal that differs between these two problems, and without loss of generality we denote it as $g_{i}$ in $\mathcal{P}$ and $g^{\prime}_{i}$ in $\mathcal{P}^{\prime}$. According to the third condition of Definition 1, signals $g_{i}$ and $g^{{}^{\prime}}_{i}$ are required to exhibit similar behaviors around the optimal solution, i.e., $g_{i}$ and $g^{{}^{\prime}}_{i}$ should converge to each other if the algorithm can ensure convergence to the optimal solution. Therefore, according to the proven convergence in Theorem 2, we can formalize this condition by stipulating the existence of a constant $C>0$ such that